OpenStudy (anonymous):

An acute angle is an angle with measure less than 90 degrees. In a right triangle, the measure of one acute angle is 27 degrees greater than two times the measure of the other acute angle. Find the measure of each acute angle. I am having trouble putting this into equation form...help:)

OpenStudy (anonymous):

Because sum of the remaining acute angles is 90 degrees

OpenStudy (mathmagician):

x+(2x+27)=180-90 3x=63 x=21 so angles are 21 and 69

OpenStudy (anonymous):

pavankumartgpk...thanks, but need the formula too:)

OpenStudy (anonymous):

The formula is sum of angles in triangle is 180 degrees

OpenStudy (anonymous):

mathmagician...you rock! Thank you!!!!

OpenStudy (mathmagician):

you are welcome :)

OpenStudy (kymber):